A solution of AgNO3 is mixed with NaCl to form a solution that is 0.10 M in AgNO3 and 0.075 M in NaCl. What will happen once these solutions are mixed? Ksp (AgCl) = 1.77
Answer #1
A solution contining AgNO3 is mixed with a solution of NaCl to form a solution that is 0.10 M in AgNO3 and 0.075 M NaCl. What will happen once these solutions are mixed? Ksp(AgCl) = 1.77×10-10.


(0.10 moles AgNO3 / L) x (1 mole Ag+ / 1 mole AgNO3) = 0.10 moles Ag+ / L = 0.10 M Ag+

(0.075 moles NaCl / L) x (1 mole Cl- / 1 mole NaCl) = 0.075 moles Cl- / L = 0.075 M Cl-

Q = [Ag+][Cl-] = (0.10)(0.075) = 0.0075 which is MUCH greater than Ksp (1.77 x 10^-10).

Since Q > Ksp, a precipitate of AgCl will form.
Answer #2
Ksp = [Ag+][Cl-] = 1.77x10^-10
Q = [Ag+][Cl-] = (0.10)(0.075) = 7.5x10^-3
Ksp < Q
Silver chloride will precipitate out of solution, leaving a saturated AgCl solution